# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ Appendix G Examples of Attacks on QKD
G.1 List of Individual Attacks
  • Large Pulse Attack [17]
@@ G.1 List of Individual Attacks
  • Photon-Number Splitting (PNS) Attack [16, 48]
    This attack is discussed in Section 3.2.1 of our work. Additional details appear in Section 2.5.1 of [30].
  • Injection-Locking Attack [22]
    This attack is discussed in Section 6.3 of our work. Additional details appear in Table 4.35 of [30].
  • Time-Shift Attack [23]
@@
    Further details can be found in Table 4.5 of [30].
  • Trojan Pony Attack [55, 32]
    This attack is initially defined in [55] as one built under a two-adversary model. In this model, an additional adversary named Fred is inside Bob’s lab, and can get classical information from Eve to affect Bob’s device in a limited way. Fred cannot communicate back to Eve. In this setup, Bob’s detector is not fully efficient, and Fred has control over when to fail Bob’s detector. When Fred knows that Eve and Bob chose mismatching bases, he drops Eve’s photon.
    A special case of this attack, given in [55], is a detector which behaves identically for pulses with a single photon and those with a high number of photons, and Bob treats double-click events as losses. In their work, [32] show that their implementation of this attack can be built using Reversed-Space.
G.2 List of Attack Families
  • Bright Illumination Attacks [21]
    The pulsed variant of this attack is analyzed in Section 7. Different variants appear in tables 4.12-4.14 and 4.17-4.20 of [30].
  • Faked States Attacks [19, 38, 20]
    This is a general family of intercept-resend attacks where Eve sends malicious signals to Bob, which cause him to get a result of her choice, as discussed in Section 3.2.3. Additional details appear in Section 2.5.2.1 of [30].
  • Fixed Apparatus Attacks [24]
@@ G.2 List of Attack Families
  • Trojan Horse Attacks [85, 18]
    Trojan Horse attacks send light pulses into Alice or Bob’s devices, and study secret aspects of their configuration from the back-scattered light. Details appear in Table 4.48 of [30]. These attacks are a general case of the Large Pulse attack, as defined by [17]. The Large Pulse attack specifically probes devices by sending a high-intensity light pulse at a specific window, whereas Trojan Horse attacks are relatively generic. Note that Lo and Chau defined a different notion of “Trojan Horse” attacks [92, 93], but it is rarely used.
  • Reversed-Space Attacks [27, 28]
@@
  • Detector Efficiency Mismatch Attacks [19, 38]
    This attack utilizes a receiver that uses several detectors in order to measure the incoming states, and their sensitivities are not completely identical. Eve uses the non-overlapping parts of the windows and sends fake states that force a specific state/basis measurement. As such, it is a special case of the Faked-States attack family. Details appear in Table 4.5 of [30].

### v3

@@ Appendix E Examples of Attacks on QKD
E.1. List of Individual Attacks
  • Large Pulse Attack (vakhitov_large_2001, )
@@ E.1. List of Individual Attacks
  • Photon-Number Splitting (PNS) Attack (PNS, ; PNS_eurocrypt, )
    This attack is discussed in Section 3.2.1 of our work. Additional details appear in Section 2.5.1 of (bsi, ).
  • Injection-Locking Attack (Pang2020, )
    This attack is discussed in Section 6.3 of our work. Additional details appear in Table 4.35 of (bsi, ).
  • Time-Shift Attack (qi_time-shift_2006, )
@@
    Further details can be found in Table 4.5 of (bsi, ).
  • This attack is initially defined in (gottesman_security_2004, ) as one built under a two-adversary model. In this model, an additional adversary named Fred is inside Bob’s lab, and can get classical information from Eve to affect Bob’s device in a limited way. Fred cannot communicate back to Eve. In this setup, Bob’s detector is not fully efficient, and Fred has control over when to fail Bob’s detector. When Fred knows that Eve and Bob chose mismatching bases, he drops Eve’s photon.
    A special case of this attack, given in (gottesman_security_2004, ), is a detector which behaves identically for pulses with a single photon and those with a high number of photons, and Bob treats double-click events as losses. In their work, (liss_practice_2020, ) show that their implementation of this attack can be built using Reversed-Space.
E.2. List of Attack Families
  • The pulsed variant of this attack (sague_blinding_pulsed_2011, ) is analyzed in Section 7. Different variants appear in tables 4.12-4.14 and 4.17-4.20 of (bsi, ).
  • This is a general family of intercept-resend attacks where Eve sends malicious signals to Bob, which cause him to get a result of her choice, as discussed in Section 3.2.3 of our work. Additional details appear in Section 2.5.2.1 of (bsi, ).
  • Fixed Apparatus Attacks (boyer_fixedapp_2014, )
@@ E.2. List of Attack Families
  • Trojan Horse Attacks (Gisin2002, ; gisin_trojan-horse_2006, )
    Trojan Horse attacks send light pulses into Alice or Bob’s devices, and study secret aspects of their configuration from the back-scattered light. Details appear in Table 4.48 of (bsi, ). These attacks are a general case of the Large Pulse attack, as defined by (vakhitov_large_2001, ). The Large Pulse attack specifically probes devices by sending a high-intensity light pulse at a specific window, whereas Trojan Horse attacks are relatively generic. Note that Lo and Chau defined a different notion of “Trojan Horse” attacks (lo_unconditional_1999, ; lo_proof_2001, ), but it is rarely used.
  • Reversed-Space Attacks (gelles_security_2012, ; gelles_reversed_2016, )
@@
  • Detector Efficiency Mismatch Attacks (makarov_faked_2005, ; makarov_effects_2006, )
    This attack utilizes a receiver that uses several detectors in order to measure the incoming states, and their sensitivities are not completely identical. Eve uses the non-overlapping parts of the windows and sends fake states that force a specific state/basis measurement. As such, it is a special case of the Faked-States attack family. Details appear in Table 4.5 of (bsi, ).